### v3

# ML4H 2025 Template: Proceedings Track

Abstract
Clinical workflows are fragmented as a patchwork of scripts and task-specific networks that often handle triage, task selection, and model deployment. These pipelines are rarely streamlined for data science pipeline, reducing efficiency and raising operational costs. Workflows also lack data-driven model identification (from imaging/tabular inputs) and standardized delivery of model outputs. In response, we present a practical, healthcare-first framework that uses a single vision–language model (VLM) in two complementary roles. First (Solution 1), the VLM acts as an aware model-card matcher that routes an incoming image to the appropriate specialist model via a three-stage workflow (modality $\rightarrow$ primary abnormality $\rightarrow$ model-card id). Checks are provided by (i) stagewise prompts that allow early exit via None/Normal/Other and (ii) a stagewise answer selector that arbitrates between the top-2 candidates at each stage, reducing the chance of an incorrect selection and aligning the workflow with clinical risk tolerance. Second (Solution 2), we fine-tune the VLM on specialty-specific datasets ensuring a single model covers multiple downstream tasks within each specialty, maintaining performance while simplifying deployment. Across gastroenterology, hematology, ophthalmology, and pathology, our single-model deployment matches or approaches specialized baselines.
Compared with pipelines composed of many task-specific agents, this approach shows that one VLM can both decide and do. It may reduce effort by data scientists, shorten monitoring, increase the transparency of model selection (with per-stage justifications), and lower integration overhead.
keywords:
Healthcare AI, Vision–Language Models, MedGemma, Model Cards, MLOps, Early termination, Confidence thresholds, Selective prediction
Figure 1: Three-stage model-card matching workflow (modality $\rightarrow$ abnormality $\rightarrow$ model card id).
Data and Code Availability
Data and code are available from the authors at reasonable request.
Institutional Review Board (IRB)
All data is publicly available and de-identified; the study is IRB-exempt.
1 Introduction
Machine learning (ML) models in healthcare face significant operational friction when moving from research prototypes to clinical practice. An estimated 90% of medical AI models never reach clinical deployment Poddar et al. (2024), and production-ready models can take weeks to months to integrate, with over half of organizations reporting 8–90 days to deploy a single model Mage AI (2025). These delays stem from several bottlenecks, and we provide two key examples here. First, to select an appropriate model design (architecture, training objective, etc.) for a given dataset, the data scientist must identify the attributes that should guide the design. In tabular data these attributes are explicit; in clinical images they are latent in the pixels—such as modality (e.g., CT) and clinical indication (e.g., acute intracranial hemorrhage)—and must first be extracted, standardized, and quality-checked before they can inform design choices. As a result, streamlined model selection for clinical imaging is often overlooked and warrants dedicated attention Willemink et al. (2020). Second, each chosen model is an operational artifact: health systems integrate, validate, monitor, and recalibrate AI solutions one by one, so developing and maintaining task-specific models multiplies the cost burden and slows adoption Brady et al. (2024).
We address these operational issues with two linked solutions: the first assists data scientists in selecting the appropriate model, and the second outlines a practical first step to shorten time-to-deployment once that selection is finalized.
Solution 1: Model-card matching
Early medical vision–language efforts aligned images and text to enable cross-modal understanding and transfer. Recently, MedGemma Sellergren et al. (2025) and UMIT Yu et al. (2025) showed that medical VLMs can approach specialized systems while maintaining generality, motivating workflows that both interpret the input and decide what to do next. As such, we adopt a medical VLM as an aware, auditable selector that reads an input image and chooses an appropriate model card via a three-stage pipeline (modality $\rightarrow$ primary abnormality $\rightarrow$ model-card id). At each stage, the VLM proposes top candidate models, and an answer selector applies a calibrated threshold that can abstain and, when appropriate, terminate early to prevent error cascades. Decisions at each stage (VLM output, any abstention, and routing to the next stage) are visible to users and logged in a consistent format, yielding transparent selection tied to model-card semantics Mitchell et al. (2019) and safe selective prediction Geifman and El-Yaniv (2019). This process accelerates development: We believe that choosing the appropriate model for a given clinical image hinges on two key pieces of information—imaging modality and disease type—details that are not always apparent to data scientists.
Solution 2: Specialty-specific deployment
Prior work shows that domain-pretrained medical encoders, when fine-tuned on target tasks, can match strong baselines across multiple datasets Zhang et al. (2022); Mei et al. (2022). In healthcare, specialty-level foundation models have also demonstrated broad coverage within a specialty—outperforming narrow, dataset-specific models across many tasks (e.g., pathology’s Prov-GigaPath achieves SOTA on 25/26 tasks across 31 tissue types) Xu et al. (2024). Operationally, fewer, broader specialty models reduce the number of artifacts a health system must validate, secure, and monitor, lowering deployment burden compared with one-model-per-dataset estates Brady et al. (2024). In line with past work, we fine-tune the same calibrated VLM at the specialty level—gastroenterology, hematology, ophthalmology, and pathology—rather than per dataset. Our goal is to match task-specific model performance while keeping everything on a single, easy-to-maintain deployment stack. Once a model card is finalized by the joint decision of the data scientist and the Solution 1 selection workflow, the data scientist can adapt the VLM to the downstream use case rather than designing a new model. This adaptation consists of tailoring the VLM prompt to the use case and eliminates the time otherwise spent building and integrating a new architecture.
Taken together, these solutions link selection and deployment in a single, calibrated workflow—a VLM that can route and execute. Compared with recent multi-agent pipelines (Shimgekar et al., 2025b, a), this minimalist design may reduce data-science workload and reduce human error. As an example, AI-driven triage has been shown to cut chest-radiograph reporting delays from 11.2 to 2.7 days in simulation while maintaining diagnostic performance Annarumma et al. (2019). Similarly, our approach may shorten the path from triage to clinical use and increase throughput without sacrificing performance.
2 Related Work
Automated matching in healthcare
Matching models based on an appropriate dataset has been explored at multiple layers of care delivery, such as patient-facing systems, specialty care, etc. Patient-facing systems (e.g., symptom checkers like Buoy Health  buoy health (2025)) narrow likely conditions and suggest next steps. Platforms such as Garner Health  Garner (2025) match patients to specialists using outcomes data. Similarly, LLM-driven controllers (HuggingGPT  Shen et al. (2023)) show how a coordinator can parse requests, consult model descriptions, and delegate to expert models. This pattern maps naturally to healthcare model cards as standardized summaries of intended use and performance Mitchell et al. (2019). Clinical imaging “orchestration” seeks to run the appropriate algorithm on the selected scan Aidoc (2024). However, prior work does not provide calibrated, stage-wise routing that is explicitly grounded in model-card semantics, with visible abstention and early termination and per-stage decision logs for audit. In the first solution, we address this gap with a three-stage router that makes every decision—VLM output, abstention, and route—user-visible and logged, cutting routing errors while preserving safety via careful prompt design and stage-specific selection logic.
Adaptation across tasks
A parallel line of work reduces packaging and integration friction for multiple models: MONAI  Project MONAI (2023) deploys standardized applications, packaging and interfaces, and self-configuring pipelines like nnU-Net  Isensee et al. (2021) that adapt architectures to new datasets with minimal manual tuning. In practice, however, maintaining one network per task is inefficient: each model has to be validated, integrated, monitored, and updated separately at the site level, duplicating effort and slowing adoption Brady et al. (2024); Boverhof et al. (2024). Multi-agent or model-zoo deployments likewise increase the number of components to secure, monitor, and recalibrate over time, compounding technical debt Sculley et al. (2015). Unlike UMIT Yu et al. (2025)—which initializes Qwen2-VL Wang et al. (2024) and follows a two-stage pipeline (full-parameter feature alignment followed by instruction fine-tuning) to reach competitive results across many datasets and modalities—there is no evidence that a single calibrated VLM can be reused across specialties within one unified deployment stack where onboarding a new task changes only the user prompt. Our second solution targets this gap: we fine-tune one medical VLM across datasets within a specialty so a single component serves multiple imaging tasks, keeping interfaces and calibration fixed while matching or approaching benchmark performance.
Table 1: Performance summary across specialties
Task External benchmark MedGemma (S–S) fine-tuned (M–S) fine-tuned
GastroenterologyPolypSet (Train: 27,048; Test: 4,719; Val: 4,214)
Polyp Classification (Acc.) 81% 61% 79%
Polyp Classification (F1 Macro) 80% 38% 77%
Polyp Classification (F1 Weighted) 81% 47% 79%
BBox Localization Acc. 78% 1%
Region Localization Acc. 32% 46%
HematologyBloodMNIST (Train: 11,959; Test: 3,421; Val: 1,712)
Cell-type Classification (Acc.) 97% 20% 98%
Cell-type Classification (F1 Macro) 8% 98%
Cell-type Classification (F1 Weighted) 10% 98%
OphthalmologyOCTMNIST (Train: 97,477; Test: 1,000; Val: 10,832)
Retinal OCT Classification (Acc.) 78% 26% 93%
Retinal OCT Classification (F1 Macro) 38% 77%
Retinal OCT Classification (F1 Weighted) 47% 79%
OphthalmologyRetinaMNIST (Train: 1,080; Test: 400; Val: 120)
Diabetic Retinopathy Grading (Acc.) 53% 55% 74%
Diabetic Retinopathy Grading (F1 Macro) 37% 59%
Diabetic Retinopathy Grading (F1 Weighted) 52% 72%
PathologyTissueMNIST (Train: 165,466; Test: 47,280; Val: 23,640)
Kidney Cortex Classification (Acc.) 70.3% 21% 70.1% 28%
Kidney Cortex Classification (F1 Macro) 5% 60%
Kidney Cortex Classification (F1 Weighted) 9% 69%
PathologyPathMNIST (Train: 89,996; Test: 7,180; Val: 10,004)
Tissue Type Classification (Acc.) 91.1% 38% 96% 21%
Tissue Type Classification (F1 Macro) 28% 94%
Tissue Type Classification (F1 Weighted) 29% 96%
PathologyBreastHIST (Train: 214,969; Test: 31,376; Val: 31,179)
Ductal Carcinoma Classification (Acc.) 84.3% 69% 87% 32%
Ductal Carcinoma Classification (F1 Macro) 66% 84%
Ductal Carcinoma Classification (F1 Weighted) 70% 87%
Figure 2: Effect of Stage 2 and Stage 3 cutoffs on the same renal ultrasound image. Top: higher $\tau_{2}{=}0.30$ with small $\tau_{3}{=}0.025$ triggers early termination (None). Bottom: lower $\tau_{2}{=}0.10$ with the same $\tau_{3}{=}0.025$ fails to halt, yielding an incorrect model-card id.
Figure 3: Two fixed-cutoff examples ($\tau_{1}{=}0.1$, $\tau_{2}{=}0.3$, $\tau_{3}{=}0.025$) with correct final selections. Top: an easy case in which the thresholds have no effect. Bottom: a case where the very small Stage 3 cutoff allows the pipeline to select the correct model code.
3 Solution 1: Model–Card Matching with Early Termination
To integrate the medical VLM as an automated model selector, we use a three-stage prompting pipeline. At each stage, the model receives a specific instruction and context to return a textual answer. The stages progressively narrow the information needed to select the appropriate model card. After each response, an answer selection step ranks all candidate answers by likelihood, retains the top two, and then decides whether to proceed to the next stage or terminate early. Figure 1 illustrates the overall workflow, and Algorithm 1 gives pseudocode for the answer-selection strategy. Below, we first outline the prerequisites and then describe the selection process for each stage.
Prerequisite 1: Backbone VLM choice
In Section 1 (Solution 1) we described a generic medical VLM; here, we adopt MedGemma. MedGemma was pretrained on a multi-modal medical mixture spanning radiology (CXR/CT/MRI), ophthalmology, dermatology, and histopathology [^1] Example datasets: MIMIC-CXR 231,483 images; EyePACS 199,258; internal dermatology 51,049; PAD-UFES-20 2,047; internal histopathology 32.55M patches; CT-US1 59,979 slices; MRI-US1 47,622 Sellergren et al. (2025). and reports competitive benchmarks relative to prior work Sellergren et al. (2025).
Prerequisite 2: Model-card repository
To match a clinical image to the correct model, we maintain a diverse repository of model cards spanning multiple modalities (e.g., CT, MRI) and AI tasks (e.g., classification, grading). Instead of storing task-specific models, each candidate is indexed by a string key—the model-card ID (a.k.a. model code; e.g., MODEL_04, which might refer to an ophthalmology model)—and two minimal descriptors that make it discoverable by the router: (i) a short task caption describing what the model does (e.g., “segments tumor regions in histopathology images”), and (ii) the modality, i.e., the scan type on which the model was trained (e.g., “breast histopathology scan”).
Stage 1: Modality identification
The goal of the first stage is to determine the imaging modality or scan type of the input image. We prompt MedGemma with a system message that casts it as a medical-imaging specialist and supply a list of admissible modalities drawn from the model-card repository. Concretely, we iterate over all model cards, extract each card’s modality field, deduplicate the values to form a unique set, and inject that set into the prompt. The prompt instructs: (1) if the image is not a medical scan, respond with “None”; (2) if it is a medical scan, choose the single most appropriate scan name from the list that matches the image content; and (3) if none of the listed modalities perfectly fits, respond with “Other.” This prompting strategy explicitly conditions the model to use the tokens “None” or “Other” as outputs in the absence of a confident match. For example, given an endoscopic image, the correct output might be “Colonoscopy”; for a non-medical photograph (e.g. a picture of a cat), the model should output “None.”
Stage 2: Primary finding detection
In the second stage, the model examines the image (knowing the modality from Stage 1) and attempts to identify the primary abnormality or finding, if any. We prompt MedGemma with an instruction such as: “Identify the single most likely disease or finding apparent in this scan. If no disease is present, respond only with: ‘Normal.’” This stage is essentially an image-based diagnosis step for a high- level finding. For example, given a colonoscopy image, the model might answer “Polyp” if one is seen; given a chest X-ray, it might answer “Normal” if nothing abnormal stands out, or a broad condition like “Pneumonia” if detected. The prompting ensures that the special token “Normal” is used to indicate the absence of any visible pathology.
Stage 3: Model-card selection
In the final stage, the model selects an appropriate card from the repository. We iterate over all cards, filter to those whose modality matches the Stage 1 choice, and pass the resulting list of (model_card_id, task_caption) pairs to the prompt. For example, if Stage 1 predicts breast histopathology scan, the candidate set might be: [MODEL_01: Classifies benign vs. malignant findings in breast histopathology slides., MODEL_04: ]
To give the model sufficient context to decide among these candidates, we include the Stage 1 modality and the Stage 2 primary abnormality as structured fields in the prompt (e.g., Patient scan details: breast histopathology scan; Primary abnormality: carcinoma). To make this a pure selection task, we mask the input image so that MedGemma acts as a language model solving a multiple-choice problem over the candidate list. The prompt instructs the model to: (i) review each candidate’s task caption; (ii) select a model only if its task explicitly matches both the identified modality and abnormality; (iii) reply None if no option fits; and (iv) when a match exists, return the model’s code (card ID) plus a one-line justification of the fit.
Answer selector and arbitration
Despite careful prompt engineering at each stage—using cues like return the best-matched, emphasizing control words (e.g., only), and explicitly listing abstention options (e.g., None)—the workflow still underweights abstentions (None, Other, Normal). This can compromise safety by increasing false selections when abstention would be preferable. Token-level analysis of secondary candidates shows that abstention terms frequently appear as the second-most likely answers. Accordingly, we shift from committing to the single top candidate to also considering the second-ranked candidate. We detail this change in the following:
Input: Candidate list $\mathcal{C}=\{(a_{i},p_{i})\}_{i=1}^{|V|}$ with $p_{1}\geq p_{2}\geq\cdots$; cutoff $\tau$; abstain set $\mathcal{A}=\{\texttt{None},\texttt{Normal},\texttt{Other}\}$.
Output: Tuple $(y,\textsf{TERMINATE})$ or $(y,\textsf{PROCEED})$.
$(a_{1},p_{1}),(a_{2},p_{2})\leftarrow\textsf{top-2 from }\mathcal{C}$; $y\leftarrow a_{1}$
if $p_{2}\geq\tau$ and $a_{2}\neq a_{1}$ then
  $y\leftarrow a_{2}$
   $\triangleright$ prefer runner-up if sufficiently likely
end if
if $y\in\mathcal{A}$ then
  return $(y,\textsf{TERMINATE})$
end if
return $(y,\textsf{PROCEED})$
Algorithm 1 Answer selector (per stage)
Selector logic
At each stage, we compute the distribution over the first token, sort the vocabulary by probability, and obtain a ranked list $\{(t_{i},p_{i})\}_{i=1}^{|\mathcal{V}|}$, where $t_{i}$ is the $i$-th ranked first token and $p_{i}$ is its probability. We then take the top two $\{(t_{i},p_{i})\}_{i=1}^{2}$ and decode full answers conditioned on each first token, yielding $\{(a_{i},p_{i})\}_{i=1}^{2}$, where $a_{i}$ is the complete answer produced when the first token is fixed to $t_{i}$. This pair of candidates is passed to a transparent answer selector that arbitrates between the two using a stage-specific cutoff $\tau$. This gives meaningful probability mass to abstentions (None/Normal/Other), which are often the runner-up.
Cutoffs in practice The purpose of per-stage cutoffs is to steer the workflow toward a more accurate outcome: either terminate early when plausible or return the best-matching model card (model code). Early termination can occur at any of the three stages—returning None or Other at Stage 1, Normal at Stage 2, or None at Stage 3—otherwise the pipeline outputs a model card id. To choose the cutoffs, we experimented on a broad set of out-of-distribution (OOD) clinical images, defined as any image outside MedGemma’s training/validation/test sets, drawn from the datasets in Section 4 and from public, license-free images Wikimedia Foundation (2025). The resulting thresholds were a low cutoff at Stage 1 ($0.10$), a moderate cutoff at Stage 2 ($0.30$), and a very small cutoff at Stage 3 ($0.025$). Figure 2 illustrates how the Stage 2 and Stage 3 thresholds can jointly flip the final outcome on the same renal ultrasound image Hansen et al. (2015). In this example, Stage 2 is incorrect (the top-two predicted abnormalities are wrong). If Stage 2 forwards its second candidate to Stage 3 (bottom panel), an incorrect model card is selected given the cutoff threshold. By contrast, if Stage 2 instead forwards the first candidate (liver cancer; top panel), the workflow abstains, returning None. Thus, the interplay between the Stage 2 candidate choice and the Stage 3 cutoff can flip the decision from a wrong selection to safe early termination. Figure 3 illustrates two fixed–cutoff cases. Bottom: the very small Stage 3 cutoff allows the pipeline to select the second-ranked answer—a plausible model-card id for a chest X-ray with cardiomegaly from the NIH Long-Tail CXR dataset Wang et al. (2017)—rather than terminate and return None. This example shows that the Stage 3 threshold was calibrated so that, when sufficiently likely, a second candidate can be prioritized over the first (abstention) option. Top: when the pipeline is highly confident at each stage—for a histopathology image with carcinoma from Cruz-Roa et al. (2014)—the cutoffs are effectively inert and do not alter the outcome.
4 Solution 2: Specialty-specific model for downstream deployment
Building on the rationale in Section 1 (Solution 2), we fine-tune the same backbone VLM adopted in Section 3MedGemma—across four specialties—gastroenterology, hematology, ophthalmology, and pathology—to simplify maintenance and deployment rather than pursue state-of-the-art performance on each task. For each specialty we compare: (a) an external task-specific benchmark (typically the published SOTA for that dataset), (b) the base MedGemma in zero-shot mode (no domain adaptation), and (c) MedGemma after specialty-specific fine-tuning (denoted “S–S” in the table). This tests whether a single specialty-tuned MedGemma can substitute for the external benchmark while keeping one deployment stack. If successful, a fixed VLM plus targeted prompt engineering and supervised fine-tuning may reduce the need to design, validate, and operate separate architectures—saving compute and shortening triage-to-deployment cycles Annarumma et al. (2019). We also evaluate multi-specialty fine-tuning; details appear below under Fine-tuning results. All results are summarized in Table 1.
Datasets
For gastroenterology, we use the dataset introduced by  Li et al. (2021), an annotated compilation of three public sources: CVC-ColonDB Bernal et al. (2012), GLRC Mesejo et al. (2016), and KUMC (80 colonoscopy video sequences from the University of Kansas Medical Center). For hematology, we use BloodMNIST Acevedo et al. (2020), organized into eight blood cell categories. For ophthalmology, we employ OCTMNIST Kermany et al. (2018) (4 retinal OCT classes) and RetinaMNIST Dataset (2020) (5 retinopathy grades). For pathology, we use TissueMNIST Ljosa et al. (2012) (8 kidney cortex cell types), PathMNIST Kather et al. (2019) (9 colorectal tissue types), and a breast cancer histopathology set following Cruz-Roa et al. (2014). Train/validation/test sizes are summarized in Table 1.
Fine-tuning results
We evaluate two approaches: single-specialty fine-tuning (training on all datasets within one specialty) and multi-specialty fine-tuning (training across multiple specialties). For both approaches, we use the same parameter-efficient fine-tuning (PEFT) method, QLoRA Dettmers et al. (2023). As shown in Table 1, single-specialty fine-tuning substantially narrows the gap to specialist benchmarks. While the base MedGemma performs modestly in zero-shot mode, targeted specialty adaptation yields competitive results across modalities and objectives. This supports the premise that one VLM can act as a generalist engine per specialty, reducing the need for separate per-task architectures and thereby saving compute and shortening triage-to-deployment cycles.
For multi-specialty fine-tuning (column “M–S fine-tuned”), we perform sequential adaptation from a pathology-tuned model to ophthalmology. Operationally, this may be preferable to single-specialty models because a single multi-specialty model can serve more inputs routed by Solution 1, reducing the number of deployed variants. As a supplementary experiment, we assess retention by measuring performance on the pathology datasets (three rows in the table). The results show that sequentially fine-tuning from pathology to ophthalmology degrades performance on pathology—an instance of catastrophic forgetting Kirkpatrick et al. (2017). This suggests that, in healthcare settings, specialty-specific models may be a better default than multi-specialty models obtained via plain sequential fine-tuning; more advanced strategies such as multi-task (joint) fine-tuning Caruana (1997) or replay buffers Rebuffi et al. (2017) may be needed to make multi-specialty models viable.
5 Discussion
We presented two linked clinical solutions: (1) a single, calibrated medical VLM that runs an aware model-card–matching workflow, routing an input image to a model-card id—or abstaining and terminating when appropriate—and (2) the same VLM used as a specialty-specific downstream model with performance competitive with task-specific models.
For (Solution 1), the workflow is designed to streamline selection for data scientists by providing: safety and trust (stagewise decisions are visible throughout the pipeline), robustness (fewer incorrect selections via the combined use of stagewise prompts and the answer-selector logic), scalability (flexibility to add stages, descriptors, or model cards, or to swap in a more mature selector), and speed (a minimalist design with fewer moving parts than recent multi-agent pipelines (Shimgekar et al., 2025b, a)).
For (Solution 2), specialty-specific models offer a practical first step toward automated deployment. We make this step concrete by shifting the effort from designing a new model to engineering a task-specific prompt that aligns the VLM’s output with the downstream objective. In practice, this process can increase throughput for data scientists and clinicians without sacrificing performance.
Limitations and future work
MedGemma is pretrained on a fixed set of modalities and tasks; accordingly, it may not generalize to imaging types or findings outside its training distribution. In short, its inference and selection decisions are tied to its training data: if a modality or pathology is unseen or underrepresented, the model may fail to recognize it and either choose an imperfect model or default to “Other/None.” This risk can be mitigated by enlarging the pretraining corpus or by adopting continual-learning methods that add new domains while preserving prior knowledge (e.g., Li and Hoiem (2017)). Another limitation is that our current model card database supports up to two tasks per model (e.g., detection followed by classification); in reality, a case might benefit from multiple models (e.g., detection, classification and grading). Extending the pipeline to return a composite plan (multiple models in sequence) is an interesting direction. It is also worth noting that while our system outputs justifications for model selection, the interpretability of the VLM’s internal reasoning remains limited (a common issue with large neural models). Techniques like chain-of-thought prompting or explicability constraints could be applied so the model provides brief, stagewise reasoning—further improving transparency Wei et al. (2022). Another design choice concerns the selection of thresholds $\tau_{1},\tau_{2},\tau_{3}$. We converged on approximate values that performed well given (i) our stage-specific prompting strategy; (ii) the diversity and structure of the model-card repository defined in Section 3; and (iii) the workflow’s subjective preference for the true final output—either None or a model-card ID—across different input images. In practice, these thresholds may require adjustment across clinical contexts. For high-accuracy applications, using the same CT of cardiomegaly in Figure 3, a workflow designer might prefer None over the closest model-card match (MODEL_13). Finally, our pipeline is across three stages because our repository lookup is primarily keyed on modality and one primary finding. As we incorporate more complex model cards (e.g., some models might handle combinations of findings or require patient metadata), the prompt design might need to extend to additional stages or a more free-form query. Fortunately, the modular nature of our approach means we can extend the prompting strategy as needed, and the answer selector concept can generalize to any number of stages by always examining the confidence of alternate responses. We outline two complementary directions for deployment: (1) automate fine-tuning end to end—so deployment proceeds with minimal data-scientist intervention—by automatically selecting the dataset linked to the model card chosen in solution 1 and automatically generating the prompts required for fine-tuning; and (2) reduce inference latency by applying pruning methods tailored to large language models (e.g., SparseGPT Frantar and Alistarh (2023)).